Order: Infuse Ceftriaxone 1 gm in 50 ml of normal saline IVPB over 30 minutes. Drip Factor: 15 gtt/ml. What is your drip rate?
Answer #1

Drip rate = gtt/min.

Gtt/min = ml/hour x drop factor/60 minutes.

Ml/hour = 50 *2 = 100 ml/hour.

Gtt/min = 100 x 15/60

Gtt/min = 25